### v2

@@ Abstract
Software development in the aerospace domain requires adhering to strict, high-quality standards. While there exist regulatory guidelines for commercial software in this domain (e.g., ARP-4754 and DO-178), these do not apply to software with deep neural network (DNN) components. Consequently, it is unclear how to allow aerospace systems to benefit from the deep learning revolution. Our work here seeks to address this challenge with a novel, output-centric approach for DNN certification. Our method employs statistical verification techniques, and has the key advantage of being able to flag specific inputs for which the DNN’s output may be unreliable — so that they may be later inspected by a human expert. To achieve this, our method conducts a statistical analysis of the DNN’s predictions for other, nearby inputs, in order to detect inconsistencies. This is in contrast to existing techniques, which typically attempt to certify the entire DNN, as opposed to individual outputs. Our method uses the DNN as a black-box, and makes no assumptions about its topology. We hope that this work constitutes another step towards integrating DNNs in safety-critical applications — especially in the aerospace domain, where high standards of quality and reliability are crucial.
Keywords:
@@
Aerospace Certification Deep Neural NetworksEnable Monitor Statistical Verification.
1 Intoduction
@@
Air transportation handles billions of passengers annually, and requires advanced solutions for addressing longstanding efficiency, environmental impact, and safety challenges [8]. The deep learning revolution, and specifically deep neural networks (DNNs), show great potential for addressing this need [28, 34]. However, although DNNs have demonstrated remarkable performance, they remain susceptible to tiny perturbations in their inputs, which may result in misclassification. This vulnerability, referred to as adversarial inputs [12, 15], could trigger a chain of events that would result in serious injury, extensive damage to valuable assets, and irreparable damage to delicate ecosystems [17]. Consequently, adversarial inputs are a key limitation that is preventing DNNs from being integrated into safety-critical systems.
@@
In the aerospace industry, certification authorities such as the European Aviation Safety Agency (EASA) and the Federal Aviation Administration (FAA) play a key role in advocating for civil aviation safety and environmental preservation. These authoritative bodies acknowledge the immense value of incorporating DNNs into safety-critical applications [8], yet also emphasize that conventional software certification guidelines, such as DO-178 [10], are not applicable to DNNs. In recognition of the inherent challenges of certifying DNNs within the aerospace domain, EASA recently released its comprehensive AI roadmap [8], which outlines seven pivotal requirements for ensuring artificial intelligence trustworthiness. Of these, robustness to adversarial conditions is mentioned as one of the fundamental prerequisites. Thus, it is evident that ensuring a DNN’s robustness to adversarial inputs will play an important role in any future certification process.
@@
During the preliminary stage of an aerospace certification process, the FAA ARP-4754 guidelines [9] call for conducting a safety analysis procedure for all aircraft systems and subsystems. For each subsystem, the safety engineering team is required to determine the permissible probability of failure, with which the design engineer then has to comply. In order to carry out such a process for a system that incorporates a DNN, the engineering team must demonstrate that the probability of a failure related to the DNN does not exceed the threshold established in the safety analysis. Unfortunately, there is currently a shortage of techniques for performing this analysis that are both sufficiently precise and scalable.
@@
Here, we seek to bridge this crucial gap by introducing a novel DNN certification approach, DNN Enable Monitor (DEM). Inspired by wireless network management concepts [20, 2], instead of trying to certify the DNN as a whole, DEM focuses on certifying only the DNN’s outputs. For each output of the DNN, DEM calculates, on-the-fly, the probability of misclassification — and if this probability is below a certain acceptable threshold, the output is considered certified. Otherwise, the output is flagged and passed on for further analysis, e.g., by a human expert. The motivation here is that well-trained DNNs, working under normal conditions, are generally correct; and so identifying the few cases where the DNN might be wrong and further analyzing only these cases can relieve a significant cognitive load off the human experts, compared to a manual approach. Further, if the automated certification is accurate enough, then the entire process becomes nearly fully automatic — and also sufficiently accurate to meet the bars specified in the safety analysis phase.
@@
Internally, DEM applies the following principle in order to determine whether an output produced by the DNN is correct or not. Consider an input $\vec{x_{0}}$, which is mapped by the DNN to label $l_{0}$. If this classification is correct, then it stands to reason that most input points very close to $\vec{x_{0}}$ would also be classified as $l_{0}$. If, however, $\vec{x_{0}}$ is an adversarial input, which can intuitively be interpreted as a “glitch” in the DNN, then other points around $\vec{x_{0}}$ would be classified differently. The purpose of DEM is to effectively distinguish between these two cases.
In order to achieve this, DEM leverages the recently proposed concept of probabilistic global categorial robustness, presented as part of the gRoMA method for statistically evaluating the global robustness of a classifer DNN, per output category [22]. The gRoMA approach is scalable, and does not require any assumptions about the DNN itself: the DNN is treated as a black-box, and the method does not depend on its activation functions, parameters, hyper-parameters, or internal topology. These advantages carry over to DEM as well. However, while gRoMA focuses strictly on measuring a DNN’s global robustness, DEM extends this approach to determine whether a particular output of the DNN should be considered certified or not.
@@
For evaluation purposes, we created a proof-of-concept implementation of DEM, and used it to evaluate VGG16 [29] and Resnet [13] DNN models. Our evaluation shows that DEM outperforms state-of-the-art methods in adversarial input detection, and that it successfully distinguishes adversarial inputs from genuine inputs — with up to 100% success rate in some categories, which is of course sufficiently high to meet the aerospace regulatory requirements.
@@
The rest of the paper is organized as follows. We begin in Sec. 2 with a description of prior work, and then present the required background on adversarial robustness in Sec. 3. We then describe our proposed method for measuring adversarial robustness in Sec. 4, followed by a description of our evaluation in Sec. 5. Finally, we summarize and discuss our results in Sec. 6.
2 Related Work
@@
Certification Guidelines. A widely accepted cornerstone in system and software certification within the aerospace domain is the “ARP-4754 — Guidelines for Development of Civil Aircraft Systems and Equipment Certification” guidelines [9]. These guidelines were authored by SAE International, a global association of aerospace engineers and technical experts. The guidelines focus on safety considerations when developing civil aircraft and their associated systems. However, these guidelines do not directly apply to DNN components.
A recent work [14] proposed an approach for assessing DNN robustness using an alternative Functional Hazard Analysis (FHA) method. This method is similar to the one required by ARP-4754, but does not quite address the requirements in the guidelines. Another recent study [11] developed a comprehensive framework of principles for DNN certification, which is aligned with ARP-4754 — but which does not include a concrete method or tool that can be used.
While ARP-4754 focuses on system level certification, the influential DO-178 guidelines for airborne systems focus on software certification [10]. These guidelines define 5 levels (A-E) for software failure conditions. Initial work has discussed certifying DNNs to level D [5, 6] or C [4, 7, 4], but to date none have reached Level A — the most severe level, required where failures may lead to fatalities, which is the case in aerospace. We hope that our approach can be used to satisfy Level A requirements in the future.
@@
One approach for drawing such conclusions is to formally verify DNNs [15, 31]. DNN verification methods are highly accurate, but suffer from limited scalability, require white-box access to the DNN in question, and pose certain limitations on the DNN’s topology and activation functions.
Another approach attempts to circumvent these limitations by certifying a DNN’s robustness with significant margins of error [35, 33, 1, 14]. This kind of attempt may be inadequate for certification in aerospace, where large error margins are usually not acceptable.
Finally, there exist statistical approaches for evaluating the probability that a specific input is being misclassified by the network (i.e., is an adversarial input)  [32, 30, 25]. However, the performance of these approaches typically falls short of the level required by regulators. DEM is also statistical at its core, and we later show that it outperforms the state of the art significantly, leading us to hope that it could be applied in aerospace. Specifically, we compare DEM to LID [23], a state-of-the-art method in this category. As we later discuss, part of the advantage of our method over LID is that DEM can be configured to be conservative, i.e., to refrain from falsely declaring that an output is certified.
3 Background
@@ A DNN $N$ is $\epsilon$-locally-robust at input point $\vec{x_{0}}$ iff
Intuitively, Definition 1 states that for input vector $\vec{x}$, the network assigns to $\vec{x}$ the same label that it assigns to $\vec{x_{0}}$, as long as the distance of $\vec{x_{0}}$ from $\vec{x}$ is at most $\epsilon$ (using the $L_{\infty}$ norm).
Probabilistic Local Robustness. Definition 1 is boolean: given $\epsilon$ and $\vec{x_{0}}$, the DNN is either robust or not robust. However, in real-world settings, and specifically in aerospace applications, systems could still be determined to be sufficiently robust if the likelihood of encountering adversarial inputs is greater than zero, but is sufficiently low. Federal agencies, for example, provide guidance that a likelihood for failure that does not exceed $10^{-9}$ (per operational hour under normal conditions) is acceptable [19]. Hence, we make an adjustment to Definition 1 that allows us to reason about a DNN’s robustness in terms of its probabilistic-local-robustness (plr[21], which is a real value that indicates a DNN’s resilience to adversarial perturbations imposed on specific inputs. More formally:
Definition 2
The $\epsilon$-probabilistic-local-robustness (PLR) score of a DNN $N$ at input point $\vec{x_{0}}$, abbreviated $\text{plr}{}{}_{\delta,\epsilon}(N,\vec{x_{0}})$, is defined as:
$\text{plr}{}{}_{\delta,\epsilon}$ $(N,\vec{x_{0}})\triangleq P_{x:\lVert\vec{x}-\vec{x_{0}}\rVert_{\infty}\leq\epsilon}[(\operatorname*{arg\,max}(N(\vec{x}))=\operatorname*{arg\,max}(N(\vec{x_{0}})))]$
Intuitively, the definition measures the probability that for an input $\vec{x_{0}}$, an input $\vec{x}$ drawn at random from the $\epsilon$-region around $\vec{x_{0}}$ will have the same label as $\vec{x_{0}}$.
Probability Global Caterorial Robustness (PGCR). Although Definition 2 is more realistic than Definition 1, it suffers from two drawbacks. The first: recent studies indicate a significant disparity in robustness among the output categories of a DNN [21, 22], which is extremely relevant to safety-critical applications. However, Definition 2 does not distinguish between output classes. The second drawback is that Definition 2 considers only local robustness, i.e., robustness around a single input point, in a potentially vast input space; whereas it may be more realistic to evaluate robustness on large, continuous chunks of the input space. To address these drawbacks, we use the following definition [22]:
Definition 3
@@
$pgcr_{\delta,\epsilon}(N,l)\triangleq P_{\vec{x_{1}},\vec{x_{2}}\in\mathbb{R}^{n}||\vec{x_{1}}-\vec{x_{2}}||_{\infty}\leq\epsilon}[|N(\vec{x_{1}})[l]-N(\vec{x_{2}})[l]|<\delta]$
@@
4 The Proposed Method
4.1 The Inference Phase
Given an input point $\vec{x}_{0}$, our objective is to quantify the reliability of the DNN’s particular prediction, $N(\vec{x}_{0})$. Our method for achieving this consists of the following steps:
@@ 4.1 The Inference Phase
  • Generate $k$ perturbations around $\vec{x}_{0}$, denoted $\vec{x}^{\prime}_{1}\ldots\vec{x}^{\prime}_{k}$, sampled uniformly at random from an $\epsilon$-region around $\vec{x}_{0}$.
  • For each sampled perturbation point $\vec{x}_{i}^{\prime}$, compute the prediction $N(\vec{x}_{i}^{\prime})$ and check whether it coincides with $N(\vec{x}_{0})$. Count the number of such matches (“number of hits”), denoted $h$.
  • If $h$ is above a certain threshold $\mathcal{T}$, certify the prediction of $N(\vec{x}_{0})$ as correct. Otherwise, flag it as suspicious.
The motivation for this approach is that, if $\vec{x}_{0}$’s prediction is genuine, the network should produce similar predictions for many points around $\vec{x}_{0}$; in which case, the number of hits will be high, and will exceed threshold $\mathcal{T}$. Conversely, if $\vec{x}_{0}$’s prediction is not genuine, i.e. $\vec{x}_{0}$ is an adversarial input for $N$, then there will be many points around $\vec{x}_{0}$ that are classified correctly — and consequently, their classification differs from that of $\vec{x}_{0}$, and the number of hits will be low.
Naturally, in order for this approach to work great care must be taken in selecting the two key parameters: $\epsilon$, which determines the size of the region around $\vec{x}_{0}$ from which perturbations are sampled, and the threshold $\mathcal{T}$ for deciding whether a prediction is correct. Intuitively, selecting too large a value for $\epsilon$ will reduce the number of hits for genuine inputs, because the region will include inputs that have true different classifications; whereas selecting too small a value will result in a high number of hits also for adversarial inputs, due to the continuous nature of DNNs. To resolve this, we present next a method for selecting these values.
4.2 Dataset Preparation and Calibration
Our approach requires an offline calibration phase, to adjust the method’s parameters to the DNN at hand. Recall that, during inference, random samples are procured from an $\epsilon$-region around the input point $\vec{x}_{0}$; and that our goal is to select $\epsilon$ in such a way that many of these samples will be classified the same way as $\vec{x}_{0}$ when it is genuine, but only a few will meet this criterion when $\vec{x}_{0}$ is adversarial. As it turns out, different $\epsilon$ values can cause dramatic differences in the number of hits; see Fig. 1 for an illustration. In order to select an appropriate value of $\epsilon$, we use an empirical approach that leverages Levy et al.’s method for computing $PGCR$ values: we test multiple potential values of $\epsilon$, and select the one that produces the best result.
Refer to captionRefer to caption
Refer to captionRefer to caption

### v3

@@ Abstract
Air transportation, a critical component of modern life, faces significant challenges concerning efficiency, environmental sustainability, and safety. Addressing these challenges necessitates innovative solutions, such as using deep neural networks (DNNs). However, although DNNs demonstrate remarkable performance, they remain susceptible to tiny perturbations in their inputs, which may result in misclassification. The vulnerability, known as adversarial inputs, may trigger a chain of events that could result in significant, or even catastrophic, failures. In this regard, adversarial inputs constitute a crucial obstacle to the integration of DNNs into safety-critical aerospace systems.
In the present work, we introduce a novel, output-centric method for certifying DNNs that addresses this challenge. This method utilizes statistical techniques to flag out specific inputs for which the DNN’s output might be unreliable so that a human expert may examine them. In contrast to existing techniques, which typically attempt to certify the entire DNN, the proposed method certifies specific outputs. Moreover, the proposed method uses the DNN as a black box and makes no assumptions about its topology.
We developed a proof-of-concept tool called DEM to demonstrate the feasibility of the proposed method. For evaluation, we tested the proposed method on a VGG-16 model, trained on the CIFAR16 dataset, showing that the proposed method achieved a high percentage of success in detecting adversarial inputs. We believe this work constitutes another step towards integrating deep neural networks in safety-critical applications — especially in the aerospace domain, where high standards of quality and reliability are crucial.
Index Terms:
@@
Aerospace Certification, Deep Neural Networks, Enable Monitor, Statistical Verification.
I Introduction
@@
Air transportation handles billions of passengers annually, and requires advanced solutions for addressing long-standing efficiency, environmental impact, and safety challenges [1]. The deep learning revolution shows significant potential for addressing these requirements, e.g., through the use of deep neural networks (DNNs) [2, 3]. However, although DNNs have demonstrated remarkable performance, they remain susceptible to tiny perturbations in their inputs, which might result in misclassification. This vulnerability, referred to as adversarial inputs [4, 5], could trigger a chain of events that might result in serious injury, extensive damage to valuable assets, or irreparable damage to delicate ecosystems [6]. Consequently, adversarial inputs pose a significant barrier to the integration of DNNs in a variety of fields, such as safety critical systems.
@@
In the aerospace industry, certification authorities such as the European Aviation Safety Agency (EASA) and the Federal Aviation Administration (FAA) play a key role in advocating civil aviation safety and environmental preservation. These authoritative bodies acknowledge the immense value of incorporating DNNs into safety-critical applications [1], yet also emphasize that conventional software certification guidelines, such as DO-178 [7], are not applicable to DNNs. In recognition of the inherent challenges of certifying DNNs within the aerospace domain, EASA recently released its comprehensive AI roadmap [1], which outlines seven pivotal requirements for ensuring artificial intelligence trustworthiness. Of these, robustness to adversarial conditions is mentioned as one of the fundamental prerequisites. Thus, it is evident that ensuring a DNN’s robustness to adversarial inputs will play a crucial role in any future certification process.
@@
During the preliminary stage of an aerospace certification process, the SAE ARP-4754 guidelines [8] call for conducting a safety analysis procedure for all aircraft systems and subsystems. For each subsystem, the safety engineering team must determine the permissible probability of failure, with which the design engineer is required to comply. In order to carry out such a process for a system that incorporates a DNN, the engineering team must demonstrate that the probability of a failure related to the DNN does not exceed the threshold established in the safety analysis. Unfortunately, there is currently a shortage of techniques for performing this analysis that are both sufficiently precise and scalable.
@@
Here, we seek to bridge this crucial gap by introducing a novel DNN certification approach, DNN Enable Monitor (DEM). Inspired by wireless network management concepts [9, 10, 11], instead of certifying the DNN as a whole, DEM focuses on certifying the concrete DNN’s outputs. For each DNN output, DEM calculates during inference time the probability of misclassification — and if this probability is below a certain acceptable threshold, the output is considered reliable. Otherwise, the output is flagged and passed on for further analysis, e.g., by a human expert. The motivation here is that well-trained DNNs, working under normal conditions, are generally correct; and so, identifying the few cases where the DNN might be wrong, and further analyzing only these cases, can relieve a significant cognitive load off the human experts, compared to a manual approach. Further, if the automated certification is accurate enough, the entire process becomes nearly fully automatic — and also sufficiently accurate to meet the bars specified in the safety analysis phase. The proposed method is in line with other semi-automatic safety analysis and optimization methods for safety-critical applications [12].
@@
In order to effectively distinguish between regular input and an adversarial input, DEM leverages the concept of probabilistic global categorial robustness, presented as part of the gRoMA method for statistically evaluating the global robustness of a classifier DNN, per output category [13]. To gain intuition, consider an input $\vec{x_{0}}$, which is mapped by the DNN to label $l_{0}$. If this classification is correct, most input points in the robustness region around $\vec{x_{0}}$ would also be classified as $l_{0}$ [13]. If, however, $\vec{x_{0}}$ is an adversarial input, which can intuitively be interpreted as a “glitch” in the DNN, then other points around $\vec{x_{0}}$ would be classified differently. DEM is used to effectively distinguish between these two cases.
@@
For evaluation purposes, we created a proof-of-concept implementation of DEM, and used it to evaluate VGG16 [14] and Resnet [15] DNN models. Our evaluation shows that DEM outperforms state-of-the-art methods in adversarial input detection, and that it successfully distinguishes adversarial inputs from genuine inputs with a very high success rate in some categories. For these categories, success rates are sufficiently high to meet the aerospace regulatory requirements.
@@
The rest of the paper is organized as follows: We begin in Sec. II with a description of prior work, and then present the required background on adversarial robustness in Sec. III. In Sec. IV we describe our proposed method for measuring adversarial robustness, followed by a description of our evaluation in Sec. V. Finally, in Sec. VI, we summarize and discuss our results.
II Related Work
@@
Certification Guidelines. A widely accepted cornerstone in system and software certification within the aerospace domain are the “ARP-4754 — Guidelines for Development of Civil Aircraft Systems and Equipment Certification” guidelines [8]. These guidelines were authored by SAE International, a global association of aerospace engineers and technical experts. The guidelines focus on safety considerations when developing civil aircraft and their associated systems. However, these guidelines do not directly apply to DNN components.
A recent work [19] proposed an approach to assessing DNN robustness through an alternative Functional Hazard Analysis (FHA) method. This method seeks to show that the likelihood that a neural network performs unexpectedly is below an acceptable threshold, defined by the user, with a confidence level of 99%. DEM, on the other hand, measures the probability of a failure condition, given a specific input. DEM meets the probabilistic goals of the ARP-4754 guidelines objectives, and could potentially be used to certify DNN-based components without further adjustments.
Another recent study [16] developed a comprehensive framework of principles for DNN certification, which is aligned with ARP-4754 — but which does not include a concrete method or tool that can be used. While ARP-4754 focuses on system level certification, the influential DO-178 guidelines for airborne systems focus on software certification [7]. The DO-178 guidelines supplement ARP-4754’s safety requirements with 5 levels (A-E) for software failure conditions. Initial work has discussed certifying DNNs to level D [18, 20] or C [21, 17], but to date none have reached Level A — the most severe level, required where failures may lead to fatalities, which is the case in aerospace. We hope that our approach can be used to satisfy Level A requirements in the future.
@@
One approach for drawing such conclusions is to formally verify DNNs [5, 22], i.e., mathematically prove that the DNN behavior adheres to a given specifications. DNN verification methods are highly accurate sound and often complete [23], but suffer from limited scalability, require white-box access to the DNN in question, and pose certain limitations on the DNN’s topology and activation functions.
Another approach attempts to circumvent these limitations by certifying a DNN’s robustness with significant margins of error [24, 25, 26, 19]. This kind of attempt may be inadequate for aerospace certification, where large error margins are usually not acceptable.
Finally, there exist statistical approaches for evaluating the probability that a specific input is being misclassified by the network (i.e., is an adversarial input) [27, 28, 29, 30, 31]. DEM is statistical at its core as well. It can be configured to be more conservative, i.e., to prefer type II error over type I error, or less, depending on the use case. A detailed comparison between DEM and a state-of-the-art technique [31] is given in section V.
III Background
@@ A DNN $N$ is $\epsilon$-locally-robust at input point $\vec{x_{0}}$ iff
Intuitively, Definition III.1 states that for input vector $\vec{x}$, the network assigns to $\vec{x}$ the same label that it assigns to $\vec{x_{0}}$, as long as the distance of $\vec{x_{0}}$ from $\vec{x}$ is at most $\epsilon$ (using the $L_{\infty}$ norm).
Probabilistic Local Robustness. Definition III.1 is Boolean: given $\epsilon$ and $\vec{x_{0}}$, the DNN is either robust or not robust. However, in real-world settings, and specifically in aerospace applications, systems could still be determined to be sufficiently robust if the likelihood of encountering adversarial inputs is greater than zero, but is sufficiently low. Federal agencies, for example, provide guidance that a likelihood that does not exceed $10^{-9}$ (per operational hour under normal conditions) for an extremely improbable failure conditions event is acceptable [34]. Consequently, prior research suggested an adjustment to Definition III.1 that allows to reason about a DNN’s robustness in terms of its probabilistic-local-robustness (plr[35], which is a real value that indicates a DNN’s resilience to adversarial perturbations imposed on specific inputs. More formally:
Definition III.2.
The $\epsilon$-probabilistic-local-robustness (PLR) score of a DNN $N$ at input point $\vec{x_{0}}$, abbreviated $\text{plr}{}{}_{\epsilon}(N,\vec{x_{0}})$, is defined as:
$\text{plr}{}{}_{\epsilon}(N,\vec{x_{0}})$ $\triangleq P_{\vec{x}:\lVert\vec{x}-\vec{x_{0}}\rVert_{\infty}\leq\epsilon}$
$\quad\left[\arg\max(N(\vec{x}))=\arg\max(N(\vec{x_{0}}))\right]$
Intuitively, the definition measures the probability that for a specific input $\vec{x_{0}}$, an input $\vec{x}$ drawn at random from the $\epsilon$-region around $\vec{x_{0}}$ will have the same label as $\vec{x_{0}}$.
Probabilistic Global Caterorial Robustness (PGCR). Although Definition III.2 is more realistic than Definition III.1, it suffers from two drawbacks. The first: recent studies indicate a significant disparity in robustness among the output categories of a DNN [35, 13], which is extremely relevant to safety-critical applications. However, Definition III.2 does not distinguish between output classes. The second drawback is that Definition III.2 considers only local robustness, i.e., robustness around a single input point, in a potentially vast input space; whereas it may be more realistic to evaluate robustness on large, continuous chunks of the input space. To address these drawbacks, we use the following definition [13]:
Definition III.3.
@@
$pgcr_{\delta,\epsilon}(N,l)$ $\triangleq P_{\vec{x_{1}},\vec{x_{2}}\in\mathbb{R}^{n},\lVert\vec{x_{1}}-\vec{x_{2}}\rVert_{\infty}\leq\epsilon}$
$\quad\left[|N(\vec{x_{1}})[l]-N(\vec{x_{2}})[l]|<\delta\right]$
@@
Hypothesis Testing. Hypothesis testing is the task of accepting the null hypothesis $H_{0}$, or rejecting it in favor of an alternative hypothesis $H_{1}$. To achieve this, we employ a statistical test $h$, which is a function of the measurements. We further select a threshold $\tau$, such that $h>\tau$ implies that $H_{0}$ is rejected in favor of $H_{1}$, whereas $h\leq\tau$ implies that $H_{0}$ is accepted.
It is common to consider the significance and power of a test in order to assess its usefulness:
  • The significance $\alpha$ is given as $\alpha=P(h>\tau\mid H_{0})$, i.e.., the probability for a false positive, leading to the rejection of $H_{0}$ although it is in fact true (type I error).
  • The power is given by $1-\beta$, where $\beta=P(h\leq\tau\mid H_{1})$ is the probability for a false negative, i.e., of wrongfully rejecting $H_{1}$ (type II error). The power can also be formulated directly as $P(h>\tau\mid H_{1})$, i.e., the probability for a true positive, or of correctly rejecting $H_{0}$, given that $H_{1}$ is true.
In general, we seek to reduce both $\alpha$ and $\beta$. Following the Neyman-Pearson lemma [36], one may maximize the power (minimize $\beta$) given $\alpha$:
$h=\frac{\mathcal{L}(H_{1}\mid\text{measurements})}{\mathcal{L}(H_{0}\mid\text{measurements})}$
where $\mathcal{L}(H\mid\text{measurements})$ is the likelihood of $H$ given the measurements, and is equal to the conditional probability $P(\text{measurements}\mid H)$. In this work, we employ such likelihood testing to decide between the hypotheses, to determine whether a given input is adversarial or not.
IV The Proposed Method
IV-A The Inference Phase
Given an input point $\vec{x}_{0}$, our objective is to quantify the reliability of the DNN’s particular prediction, $N(\vec{x}_{0})$; i.e., to accept or reject the $H_{0}$ hypothesis that the input is adversarial. Our method for achieving this consists of the following steps:
@@ IV-A The Inference Phase
  • Generate $k$ random perturbations around $\vec{x}_{0}$, denoted $\vec{x}^{\prime}_{1}\ldots\vec{x}^{\prime}_{k}$, sampled uniformly at random from an $\epsilon$-region around $\vec{x}_{0}$.
  • For each sampled perturbation point $\vec{x}_{i}^{\prime}$, compute the prediction $N(\vec{x}_{i}^{\prime})$ and check whether it coincides with $N(\vec{x}_{0})$. Count the number of such matches (“number of hits”), denoted as $h$.
  • If $h$ is above a certain threshold $\mathcal{T}$, certify the prediction of $N(\vec{x}_{0})$ as correct, i.e., reject the $H_{0}$ hypothesis. Otherwise, flag it as suspicious.
The motivation for this approach is based on III.3. Intuitively, for a genuine prediction over the input $\vec{x}_{0}$, the network would produce stable predictions in the $\epsilon$-region around $\vec{x}_{0}$; in such a case, the number of hits will be high and will exceed the threshold $\mathcal{T}$. Conversely, if $\vec{x}_{0}$’s prediction is not genuine, i.e. $\vec{x}_{0}$ is an adversarial input for $N$, then the network would produce different predictions in the $\epsilon$-region around $\vec{x}_{0}$, and the number of hits will be lower.
Naturally, in order for this approach to work, care must be taken in determining the two key parameters: $\epsilon$, which determines the size of the region around $\vec{x}_{0}$ from which perturbations are sampled, and the threshold $\mathcal{T}$ for deciding whether a prediction is correct. Below we present a method for selecting these values.
IV-B Dataset Preparation and Calibration
Our approach requires an offline calibration phase, to adjust the method’s parameters to the DNN at hand. Recall that, during inference, random samples are procured from an $\epsilon$-region around the input point $\vec{x}_{0}$; and that our goal is to select $\epsilon$ in such a way that many of these samples will be classified the same way as $\vec{x}_{0}$ when it is genuine, but only a few will meet this criterion when $\vec{x}_{0}$ is adversarial. As it turns out, different $\epsilon$ values can cause dramatic differences in the number of hits; see Fig. 1 for an illustration. In order to select an appropriate value of $\epsilon$, we use an empirical approach that leverages Levy et al.’s method [13] for computing PGCR values: we test multiple potential $\epsilon$ values, and select the one that produces the most accurate result.